6. Prove the following relation: ] = -Ꮴ2 ᏧᏙᎢ .
Answer #1

The given expressianis ar ($), (SO, = -v)($()), an () = =v*(**) (*). [C+ -] Naw by definitian, v= f (TP) dva (SV) JT + CephedH = Tidst vode ($#), - T. (697) + V-in) Again fram Maxwells relatian, for = vedpasidt STUP Naw by Eulers reciprocal theramNaw meblitting (85/5p), in eq-ir) *), = 7.65%),+v (3 4%87) - V-76), proved.